OpenStudy (anonymous):

FAN AND MEDAL! PLEASE HELP ME! THIS IS THE LAST QUESTION I NEED TO ANSWER AND I CAN'T FIGURE IT OUT RIGHT ): The two triangles below are similar. What is the similarity ratio of ∆ABC to ∆XYZ? Select one: a. 2:3 b. 2:1 c. 3:2 d. 1:2

563blackghost (563blackghost):

I would need to see a picture of the two triangles in order to get the ratio...

563blackghost (563blackghost):

says I cant access....sorry about that

OpenStudy (anonymous):

Does that work?

563blackghost (563blackghost):

Yup!

OpenStudy (anonymous):

Please help ): I have 10 minutes to get this haha..

OpenStudy (anonymous):

Its probably really simple, i just keep getting the answer wrong..

563blackghost (563blackghost):

15 could become either 10 or 30... So lets first try with 10 \[\frac{ 15 }{ 10 } \div \frac{ 5 }{ 5 } = \frac{ 3 }{ 2 } = 3:2\] Now we simplify to its lowest and we get the ratio 3:2 now lets try 30.... \[\frac{ 15 }{ 30 } \div \frac{ 15 }{ 15 }=\frac{ 1 }{ 2 }=3:2\] Now this one simplifies dow to the ratio 1:2..... Now we have our ratio but we need to remeber that we have to do it according to corresponding sides so 15 corresponds to which side?...

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

so it's 3:2?

563blackghost (563blackghost):

Correct! So the ratio is 3:2 which is C

OpenStudy (anonymous):

Yay! you're a life saver. Thank you so much :)

563blackghost (563blackghost):

Np! Glad to be of help ^^